OpenStudy (anonymous):

for classes of 279 students, find the mean and standard deviation for the number born on 4th of July. ignore leap year

OpenStudy (kropot72):

The probability of any student being born on 4th of July is 1/365. The binomial distribution applies therefore: \[mean=np=279\times \frac{1}{365}=you\ can\ calculate\] \[standard\ deviation=\sqrt{np(1-p})=\sqrt{279\times\frac{1}{365}\times\frac{364}{365}}\]